OpenStudy (anonymous):

Evaluate the integral. x(x-2) dx from 0 to 4

OpenStudy (tkhunny):

Hint: \[\int x\cdot(x-2)\;dx\;=\;\int x^{2} - 2x\;dx\]

OpenStudy (anonymous):

Thank you. Now I got it.